# Constraining anti-baryonic dark matter through correlated nucleon decay signatures

### v1

@@ Abstract
Baryon number violation in the visible sector induced by anti-baryonic dark matter provides a viable mechanism for low-scale baryogenesis. Two of the most sensitive probes of this scenario are neutron decay processes such as $n\to\text{invisible}$ and $n\to\pi^{0}+\text{invisible}$. In this work, we discuss the generation of di-nucleon decay processes such as $nn\to\bar{\nu}\bar{\nu}$ and $nn\to\pi^{0}\pi^{0}$ at one-loop, arising from the operators responsible for induced nucleon decays. While nucleon decay rates in this model depend on the local dark matter density, di-nucleon decay processes do not, providing a complementary probe of the new physics. We thus use bounds on both nucleon and di-nucleon decays to constrain the mass and terrestrial density of anti-baryonic dark matter.
@@ I Introduction
In recent years, low-scale dark matter-induced baryon number violation Kaplan:2009ag ; Haba:2010bm ; Davoudiasl:2010am ; Davoudiasl:2011fj ; petraki2013review ; Huang:2013xfa ; Arun:2022eqs ; Akshay:2022vkb has emerged as a compelling framework to bridge this gap against the backdrop of baryogenesis mechanisms involving high-scale interactions Davidson:2008bu ; Grossman:2018rdg . Such frameworks often predict observable nucleon decay channels with new physics interactions of dark matter charged with anti-baryon number that allow for baryon number violation at experimentally accessible scales, unlike the severely restricted baryon number violation at Grand Unification scales. In these ’asymmetric’ dark matter models, proton and neutron decay can be induced typically via non-renormalizable operators like,
@@
$\mathcal{O}_{udd\bar{\nu}}=\frac{1}{\Lambda_{udd\rightarrow\bar{\nu}}^{4}}\hskip 2.84544ptu_{r}d_{r}Q_{l}L_{l}\Phi_{i}^{\dagger}\Phi_{j}\ ,$ $\mathcal{O}_{udd\pi^{0}}=\frac{1}{\Lambda_{udd\rightarrow\pi^{0}}^{3}}\hskip 2.84544ptu_{r}d_{r}d_{r}\Psi_{i}\Phi_{j}^{\dagger}\ ,$ (1)
where $\Phi_{i}$ and $\Psi_{i}$ are dark sector scalar and fermionic fields, and $Q_{l},L_{l}$ are left-handed SM doublets, while $u_{r},d_{r}$ are right-handed SM singlets. The dark sector fields are Standard Model gauge singlets, however, they are charged under the baryon number, $U(1)_{B}$, carrying anti-baryonic charge. Note that, above, we have considered only the minimal set of operators that preserve (violate) $B-L$ in the visible sector on the left (right). These operators induce decay channels like,
@@
[…]
$\mathcal{O}_{n\nu}=\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}}\hskip 2.84544ptn\nu_{l}\Phi_{1}^{\dagger}\Phi_{2}\ ,$ $\mathcal{O}_{n\pi^{0}}=\frac{1}{\Lambda_{n\rightarrow\pi^{0}}}\hskip 2.84544ptn\pi^{0}\Psi\Phi^{\dagger}\ .$ (3)
@@
Given that the asymmetric dark matter models ingeniously bridge the gap between dark sector and baryon number violation in the visible sector, and prospects of discovering $\Delta B=2$ in upcoming experiments, it is sensible to ask whether there exists correlation of dark matter induced visible sector baryon number violation processes with $\Delta B=2$ decays. Any confirmation of neutron-antineutron oscillation or di-nucleon decay with back-to-back Cherenkov radiation would indicate spontaneous breaking baryon number by two units in the dark sector Arun:2022eqs . Specially in asymmetric dark matter models, this generates di-nucleon decays at one-loop order from dimension-5 hadronic operators in Eq. (3).
[…]
@@
   $n\to\text{invisible}$    $5.8\times 10^{29}$ KamLAND:2005pen ; JUNO:2024pur $3.60\times 10^{-62}$
@@
   $nn\to\bar{\nu}\bar{\nu}$ $1.4\times 10^{30}$ KamLAND:2005pen ; JUNO:2024pur $1.49\times 10^{-62}$
@@
Table 1: Lifetimes and decay widths for the relevant processes.
The article is presented as follows. In section II, we discuss the dark sector interactions with visible sector baryon number violating operators, emphasizing on two different dark matter models that lead to $n\to\bar{\nu}$, which keeps the $B-L$ preserved, and $n\to\pi^{0}$, which violates $B$. In section III, we develop models for spontaneous breaking of the dark sector baryon charge and derive the di-nucleon decay operators that are generated at one-loop from the induced baryon number violating operators. We derive, in section IV, the cross-sections for the induced baryon number violating processes, and correlate them with the matching provided at one-loop with the di-nucleon decay and compare them with the experimental values provided in Table. (1). We then rewrite these results as a function of dark matter density and mass of dark matter and provide limits that satisfy the di-nucleon processes and relic density bound. In section V we summarize our findings.
@@ II Dark matter induced nucleon decay
In this section, we briefly recap the relevant concepts Davoudiasl:2010am ; Davoudiasl:2011fj ; Huang:2013xfa with regard to induced baryon number violation by a dark matter carrying anti-baryon number. The stability of the particles involved in the processes described in Eq. (3) are ensured if we assume the condition,
$|m_{\Phi_{1}}-m_{\Phi_{2}}|$ $<$ $m_{p}+m_{e}<m_{\Phi_{1}}+m_{\Phi_{2}}$
$|m_{\Phi}-m_{\Psi}|$ $<$ $m_{p}+m_{\pi}<m_{\Phi}+m_{\Psi}\ ,$ (4)
where these are the masses of proton, electron and the dark matter candidates. Though both the proton and the neutron can take part in these processes, since their limits are comparable, we choose to work with neutrons. The same analysis can be extended for protons without loss of generality.
With these restrictions in place, the decay processes in Eq. (3) generate the effective rate of the induced neutron decay,
@@
where, $\rho_{DM}$ is the density of the dark matter, $M_{DM}$ is the mass of dark matter and $(\sigma v)$ is the thermal cross-section for the particular nucleon decay. In what follows, we will consider two scenarios and compute the induced thermal cross-section $(\sigma v)_{IND}$.
@@
To start, let’s consider a neutron decay scenario induced by two scalar dark matter (DM) fields $\Phi_{1}$ and $\Phi_{2}$, for the decay channel $n+\Phi_{1}\rightarrow\bar{\nu}+\Phi_{2}$. The quark-level Lagrangian for the decay is defined as
@@
where the charges of the fields are assigned in Table. (2). Both the dark sector fields considered here are scalars and are nearly degenerate in mass for stability reasons.
$SU(3)_{c}$ $SU(2)_{W}$ $U(1)_{Y}$ $U(1)_{B}$ $U(1)_{L}$ $U(1)_{D}$
$(u_{r},d_{r})$ 3 1 (4/3, -2/3) 1/3 0 0
$Q_{l}$ 3 2 1/3 1/3 0 0
$L_{l}$ 1 2 1 0 1 0
$\Phi_{1}$ 1 1 0 1 0 1
$\Phi_{2}$ 1 1 0 0 -1 1
@@
$\mathcal{L}_{n\bar{\nu}}=\hskip 1.42271pt\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}}\hskip 2.84544ptn\ \nu_{l}\hskip 2.84544pt\Phi_{1}^{\dagger}\Phi_{2}\ ,$ (7)
where, $\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}}=\frac{\alpha}{\Lambda_{udd\rightarrow\bar{\nu}}^{4}}$, is the scale of new physics and $\alpha=0.015\hskip 2.84544pt\mathsf{GeV}^{3}$ gives the hadronic matrix element. Here, we follow the conventions given in Claudson:1981gh . This operator generates the decay as shown in Fig. (1).
@@
An alternative scenario, in which the $B-L$ symmetry is violated, involves considering neutron decay into a pion within the visible sector. The Lagrangian describing this process is given by Davoudiasl:2010am ; Davoudiasl:2011fj ,
$\mathcal{L}_{udd\pi}=\frac{1}{\Lambda_{udd\pi}^{3}}\Phi^{\dagger}\Psi_{R}d_{R}d_{R}u_{R}+h.c.\ ,$ (9)
where the charges of the fields are given in Table. (3). Here, we consider a fermionic dark sector field $\Psi$ and a scalar dark sector field $\Phi$, which are nearly degenerate in mass.
$SU(3)_{c}$ $SU(2)_{W}$ $U(1)_{Y}$ $U(1)_{B}$ $U(1)_{D}$
$(u_{r},d_{r})$ 3 1 (4/3, -2/3) 1/3 0
$\Psi$ 1 1 0 $-\frac{1}{2}$ 1
$\Phi$ 1 1 0 $\frac{1}{2}$ 1
@@
where the Feynman diagram for the decay is shown in Fig. (2). The hadronic operator is matched with the quark-level operator to get,
@@
where, $F_{\pi}=93\hskip 2.84544pt\mathsf{MeV}$ is the pion decay constant.
Figure 2: The decay of $n\rightarrow\pi^{0}$
Therefore, the thermal cross-section for the process $n+\Phi\to\pi^{0}+\Psi$ becomes,
@@
The decay width given in Eq. (5) can now be computed by using Eq. (8) and Eq. (12) as,
@@
Since these results depend on the anti-baryonic dark matter density and the scale of the new physics, an independent probe is highly appreciated.
@@ III Spontaneous breaking of dark baryon number and generation of di-nucleon decay
[…]
Upon spontaneously breaking the $U(1)_{B}$ and $U(1)_{L}$ symmetries through vacuum expectation values of $v_{S_{1}}=\langle S_{1}\rangle$ and $v_{S_{2}}=\langle S_{2}\rangle$. As a result, di-nucleon decay, $nn\to\bar{\nu}\bar{\nu}$ emerge at one-loop from Eq. (7) and Eq. (14) as shown in Fig. (3).
@@
On the other hand, the effective Lagrangian for di-nucleon decay of neutrons to anti-neutrinos, at the hadron level can be written as,
$\mathcal{L}_{nn\bar{\nu}\bar{\nu}}=\frac{1}{\Lambda_{nn\rightarrow\bar{\nu}\bar{\nu}}^{2}}(n\nu_{l})\hskip 2.84544pt(n\nu_{l})\ ,$ (15)
where $\Lambda_{nn\rightarrow\bar{\nu}\bar{\nu}}$ denotes the characteristic scale of the new physics responsible for the baryon number violating process. The one-loop process in Fig. (3) generates the operator in Eq. (15) from the nucleon decay operator in Eq. (7), and their matching becomes,
$\frac{1}{\Lambda_{nn\rightarrow\bar{\nu}\bar{\nu}}^{2}}=\frac{1}{16\pi^{2}}\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}^{2}}\frac{(\lambda_{1}v_{S_{1}})^{2}(\lambda_{2}v_{S_{2}})^{2}}{M_{s}^{4}}=\frac{1}{16\pi^{2}}\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}^{2}}\frac{m_{\Phi}^{4}}{M_{s}^{4}}\ .$ (16)
In the above relation, $M_{s}$ denotes the loop momentum and is taken to be $M_{s}=0.1\ \Lambda_{n\rightarrow\bar{\nu}}$ for calculations. Using the decay width for this process,
$\Gamma_{nn}\sim\frac{m_{n}^{5}}{\Lambda_{nn\rightarrow\bar{\nu}\bar{\nu}}^{4}}$ (17)
and the experimental lower limit on the lifetime for di-nucleon decay given in Table. (1), the constraint on the new physics scale becomes,
$\frac{1}{\Lambda_{nn\rightarrow\bar{\nu}\bar{\nu}}^{2}}<1.43\times 10^{-31}\hskip 2.84544pt\mathsf{GeV}^{-2}$ (18)
Considering, $\lambda_{1}v_{S_{1}}=\lambda_{2}v_{S_{2}}=m_{\Phi}=10\hskip 2.84544pt\mathsf{GeV}$, the independent limit on $\Lambda_{n\rightarrow\bar{\nu}}$ can be derived as,
$\Lambda_{n\rightarrow\bar{\nu}}>1.28\times 10^{6}\hskip 2.84544pt\mathsf{GeV}$ (19)
@@
Unlike in the previous scenario, here we consider two dark sector scalar fields $S$ and $\Delta$ which are SM singlets but carry baryon number. Their interaction with the dark matter fields then become,
$\mathcal{L}_{int}=\lambda_{1}m_{\phi}S\Phi^{2}+\lambda_{2}\Delta\Psi^{c}\Psi+c.c\ ,$ (20)
As before, di-nucleon decay $nn\to\pi^{0}\pi^{0}$ is generated at one-loop from the neutron decay operators in Eq. (10), as shown in Fig. (4), following the spontaneous breaking of $U(1)_{B}$ symmetry through the vacuum expectation values $v_{S}=\langle S\rangle$ and $v_{\Delta}=\langle\Delta\rangle$.
@@
The hadron-level effective Lagrangian for the di-neutron decay ($nn\to\pi^{0}\pi^{0}$) is given by,

### v2

@@ Abstract
Baryon number violation in the visible sector induced by anti-baryonic dark matter provides a viable mechanism for low-scale baryogenesis. Two of the most sensitive probes of this scenario are neutron decay processes such as $n\to\bar{\nu}+\text{invisible}$ and $n\to\pi^{0}+\text{invisible}$. In this work, we discuss the possible spontaneous breaking of baryon symmetry in the dark sector and the generation of di-nucleon decay processes such as $nn\to\bar{\nu}\bar{\nu}$ and $nn\to\pi^{0}\pi^{0}$ at one-loop, arising from the operators responsible for induced nucleon decays. While the induced nucleon decay rates in this model depend on the dark matter density, di-nucleon decay processes do not, providing a complementary probe of the new physics. We thus use nucleon and di-nucleon decay bounds to constrain the local density and mass of the anti-baryonic dark matter.
@@ I Introduction
In recent years, low-scale dark matter-induced baryon number violation Kaplan:2009ag ; Haba:2010bm ; Davoudiasl:2010am ; Davoudiasl:2011fj ; petraki2013review ; Huang:2013xfa ; Arun:2022eqs ; Akshay:2022vkb has emerged as a compelling framework to bridge this gap against the backdrop of baryogenesis mechanisms involving high-scale interactions Davidson:2008bu ; Grossman:2018rdg . Such frameworks often predict observable induced nucleon decay channels with new physics interactions of dark matter charged with anti-baryon number that allow for baryon number violation at experimentally accessible scales, unlike the severely restricted baryon number violation at Grand Unification scales. In these "asymmetric" dark matter models, proton and neutron decay can be induced typically via non-renormalizable operators like,
@@
$\mathcal{O}_{udd\nu}=\frac{1}{\Lambda_{udd\rightarrow\nu}^{4}}\hskip 2.84544ptu_{r}d_{r}Q_{l}L_{l}\Phi_{i}^{\dagger}\Phi_{j}\ ,$ $\mathcal{O}_{udd\pi^{0}}=\frac{1}{\Lambda_{udd\rightarrow\pi^{0}}^{3}}\hskip 2.84544ptu_{r}d_{r}d_{r}\Psi_{i}\Phi_{j}^{\dagger}\ ,$ (1)
where $\Phi_{i}$ and $\Psi_{i}$ are dark sector scalar and fermionic fields, and $Q_{l},L_{l}$ are left-handed SM doublets, while $u_{r},d_{r}$ are right-handed SM singlets. The dark sector fields are Standard Model gauge singlets, however, they are charged under the baryon number, $U(1)_{B}$, carrying anti-baryonic charge and/or under the Lepton number, $U(1)_{L}$, carrying anti-leptonic charge. In this work, the $U(1)_{B}$ and $U(1)_{L}$ groups are assumed to be gauged, and the associated anomalies are canceled either by a minimal expansion of the fermionic sector FileviezPerez:2010gw ; FileviezPerez:2024fzc or by Green-Schwarz mechanism by including axion interactions Green:1984sg ; Anastasopoulos:2006cz ; Coriano:2007fw . Note that, above, we have considered only the minimal set of operators that preserve (violate) BLB-L in the visible sector on the left (right).
$\mathcal{O}_{pe^{+}}=\frac{1}{\Lambda_{p\rightarrow e^{+}}}\hskip 2.84544ptpe_{l}\Phi_{1}^{\dagger}\Phi_{2}\ ,$ $\mathcal{O}_{p\pi^{+}}=\frac{1}{\Lambda_{p\rightarrow\pi^{+}}}\hskip 2.84544ptp\pi^{+}\Psi\Phi^{\dagger}\ ,$
$\mathcal{O}_{n\nu}=\frac{1}{\Lambda_{n\rightarrow\nu}}\hskip 2.84544ptn\nu_{l}\Phi_{1}^{\dagger}\Phi_{2}\ ,$ $\mathcal{O}_{n\pi^{0}}=\frac{1}{\Lambda_{n\rightarrow\pi^{0}}}\hskip 2.84544ptn\pi^{0}\Psi\Phi^{\dagger}\ .$ (2)
The stability of the particles involved in these processes are ensured if we assume the condition,
$|m_{\Phi_{1}}-m_{\Phi_{2}}|$ $<$ $m_{p}+m_{e}<m_{\Phi_{1}}+m_{\Phi_{2}}$
$|m_{\Phi}-m_{\Psi}|$ $<$ $m_{p}+m_{\pi}<m_{\Phi}+m_{\Psi}\ ,$ (3)
where these are the masses of proton, electron, pion and the dark matter candidates. Similarly for neutrons also. These operators generate ‘induced’ decay channels like,
@@
Given that asymmetric dark matter models naturally couple the dark sector to baryon number violation in the visible sector, and that next-generation searches are sensitive to $\Delta B=2$ processes, it is well motivated to ask whether dark sector induced baryon number violation in the visible sector correlates with $\Delta B=2$ decays. Any confirmed signal of neutron-antineutron oscillations or of di-nucleon decay with back-to-back Cherenkov rings would point to spontaneous breaking of baryon number by two units in the dark sector Arun:2022eqs . In this framework, the same spurion of $\Delta B=1$ breaking typically generates di-nucleon decays at one loop via the dimension-5 hadronic operators in Eq. (2), thereby linking the $\Delta B=2$ nuclear decay rates to the underlying symmetry-breaking parameters of the dark sector.
In this work, we provide a model for such breaking of baryon symmetry in the asymmetric dark sector and examine the implications of dark matter interactions on baryon number violation, focusing on the processes $n\to\bar{\nu}+\text{invisible}$, $\to\pi^{0}+\text{invisible}, nn$$n\to\bar{\nu}\bar{\nu}$ and $nn\to\pi^{0}\pi^{0}$. The experimental data relevant for this study are given in Table. (1). Using measurements of these nucleon and di-nucleon processes, we find that the mass and density of anti-baryonic dark matter at terrestrial experiments is constrained.
@@
$n\to\text{invisible}$ $5.8\times 10^{29}$ KamLAND:2005pen $3.60\times 10^{-62}$
@@
$nn\to\bar{\nu}\bar{\nu}$ $1.4\times 10^{30}$ KamLAND:2005pen $1.49\times 10^{-62}$
@@
Table 1: Limit on lifetimes and decay widths for the relevant processes from experiments.
The article is presented as follows. In section II, we discuss the dark sector interactions with visible sector baryon number violating operators, emphasizing on two different dark matter models that lead to $n\to\bar{\nu}$, which keeps the BLB-L preserved, and $n\to\pi^{0}$, which violates $B$. In section III, we develop models for spontaneous breaking of the dark sector baryon charge and derive the di-nucleon decay operators that are generated at one-loop from the induced baryon number violating operators. We derive, in section IV, the cross-sections for the induced baryon number violating processes, and correlate them with the matching provided at one-loop with the di-nucleon decay and compare them with the experimental values provided in Table. (1). We then rewrite these results as a function of dark matter density and mass of dark matter and provide limits that satisfy the di-nucleon processes and local density bound. In section V we summarize our findings.
@@ II Dark matter induced nucleon decay
In this section, we briefly recap the relevant concepts Davoudiasl:2010am ; Davoudiasl:2011fj ; Huang:2013xfa concerning induced baryon number violation by dark matter carrying anti-baryon number. Although both protons and neutrons can participate in these processes, we restrict our analysis to neutrons. The induced nucleon decay rate is given by,
@@
where, $\rho_{DM}$ is the density of the dark matter, $M_{DM}$ is the mass of dark matter and $(\sigma v)$ is the thermal cross-section for the particular induced nucleon decay generated by the operators in Eq. (2). In the following, we consider two scenarios and compute the corresponding induced thermal cross-section, $(\sigma v)_{IND}$.
@@
To begin, we consider a neutron decay scenario induced by two scalar dark matter (DM) fields, $\Phi_{1}$ and $\Phi_{2}$, corresponding to the channel $n+\Phi_{1}\rightarrow\bar{\nu}+\Phi_{2}$. The quark-level Lagrangian governing this decay is given by,
@@
where the charges of the fields, under the gauge groups, are assigned in Table. (2). As mentioned in the introduction, the anomalies associated with gauging $U(1)_{B}$ and $U(1)_{L}$ are assumed to be canceled FileviezPerez:2010gw ; FileviezPerez:2024fzc ; Green:1984sg ; Anastasopoulos:2006cz ; Coriano:2007fw . Both the dark sector fields considered here are scalars and are nearly degenerate in mass for stability reasons.
$SU(3)_{c}$ $SU(2)_{W}$ $U(1)_{Y}$ $U(1)_{B}$ $U(1)_{L}$
$(u_{r},d_{r})$ 3 1 (4/3, -2/3) 1/3 0
$Q_{l}$ 3 2 1/3 1/3 0
$L_{l}$ 1 2 1 0 1
$\Phi_{1}$ 1 1 0 1 0
$\Phi_{2}$ 1 1 0 0 -1
@@
$\mathcal{L}_{n\bar{\nu}}=\hskip 1.42271pt\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}}\hskip 2.84544ptn\ \bar{\nu}_{l}\hskip 2.84544pt\Phi_{1}^{\dagger}\Phi_{2}\ ,$ (7)
where, $\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}}=\frac{\alpha}{\Lambda_{udd\rightarrow\bar{\nu}}^{4}}$, is the scale of new physics and $\alpha=0.015\hskip 2.84544pt\mathsf{GeV}^{3}$ gives the hadronic matrix element Claudson:1981gh . This operator generates the induced decay process illustrated in Fig. (1).
@@
An alternative scenario, in which the BLB-L symmetry is violated, involves neutron decay into a pion within the visible sector. The corresponding Lagrangian is given by Davoudiasl:2010am ; Davoudiasl:2011fj ,
$\mathcal{L}_{udd\pi}=\frac{1}{\Lambda_{udd\pi}^{3}}\Phi^{\dagger}\Psi d_{R}d_{R}u_{R}+h.c.\ ,$ (9)
where the charges of the fields are given in Table. (3). Like before, the anomalies associated with the $U(1)_{B}$ are assumed to be canceled by extending the fermionic sector or via axion interactions. Here, we consider a fermionic dark sector field $\Psi$ and a scalar dark sector field $\Phi$, which are nearly degenerate in mass.
$SU(3)_{c}$ $SU(2)_{W}$ $U(1)_{Y}$ $U(1)_{B}$
$(u_{r},d_{r})$ 3 1 (4/3, -2/3) 1/3
$\Psi$ 1 1 0 $\frac{1}{2}$
$\Phi$ 1 1 0 $\frac{1}{2}$
@@
where the Feynman diagram for the induced decay is shown in Fig. (2). The hadronic operator is matched with the quark-level operator to get,
@@
where, $F_{\pi}=93\hskip 2.84544pt\mathsf{MeV}$ is the pion decay constant. And, the thermal cross-section for the process $n+\Phi\to\pi^{0}+\Psi$ becomes,
@@
Figure 2: The decay of $n\rightarrow\pi^{0}$
The induced decay width given in Eq. (5) can now be computed by using Eq. (8) and Eq. (12) as,
@@
The limits on the new physics scale can be obtained by combining the experimental constraints on the nucleon decay widths listed in Table.(1) with the above expression for the induced decay width. In particular, using the current experimental bounds from Table.(1), the resulting constraints on the scales associated with the induced neutron decay channels ($n\rightarrow\bar{\nu}$ and $n\rightarrow\pi$), for a benchmark dark matter density of $\rho_{DM}=0.3\ \mathsf{GeV/cm^{3}}$ and dark matter mass $M_{DM}=2\ \mathsf{GeV}$, are summarized in Table. (4). Since these constraints rely sensitively on both the local anti-baryonic dark matter density and the assumed scale of new physics, complementary and independent probes are therefore highly desirable.
   Models   Decay Reaction   $\Lambda_{n}$   $\sigma v$
$(\mathsf{GeV})$ $(\mathsf{GeV^{-2}})$
Case- 1 $n\rightarrow\bar{\nu}$ $1.03\times 10^{9}$ $3.12\times 10^{-20}$
Case- 2 $n\rightarrow\pi^{0}$   $7.31\times 10^{10}$   $1.63\times 10^{-23}$
Table 4: Limit on induced neutron decay scales.
@@ III Spontaneous breaking of dark baryon number and generation of di-nucleon decay
$\mathcal{L}_{int}=\frac{1}{2}\Big(\lambda_{1}m_{\phi_{1}}S_{1}\Phi_{1}^{2}+\lambda_{2}m_{\phi_{2}}S_{2}\Phi_{2}^{2}+\text{c.c}\Big)+m_{\phi_{1}}^{2}\Phi_{1}^{\dagger}\Phi_{1}+m_{\phi_{2}}^{2}\Phi_{2}^{\dagger}\Phi_{2}\ .$ (14)
Upon spontaneously breaking the $U(1)_{B}$ and $U(1)_{L}$ symmetries through vacuum expectation values of $v_{S_{1}}=\langle S_{1}\rangle$ and $v_{S_{2}}=\langle S_{2}\rangle$, di-nucleon decay, $nn\to\bar{\nu}\bar{\nu}$ arise at one-loop from Eq. (7) and Eq. (14) as illustrated in Fig. (3). The associated Goldstone bosons, in the unitary gauge, become non-dynamical or gets "eaten away" by the gauge fields of $U(1)_{B}$ and $U(1)_{L}$. Thus, there are no massless modes in the energy spectrum of the model.
@@
The effective Lagrangian for the di-nucleon decay process at the hadron level can be written as,
$\mathcal{L}_{nn\bar{\nu}\bar{\nu}}=\frac{1}{\Lambda_{nn\rightarrow\bar{\nu}\bar{\nu}}^{2}}(n\bar{\nu}_{l})\hskip 2.84544pt(n\bar{\nu}_{l})\ ,$ (15)
where $\Lambda_{nn\rightarrow\bar{\nu}\bar{\nu}}$ denotes the characteristic scale of the new physics responsible for the baryon number violating process. The one-loop process shown in Fig. (3) generates the operator in Eq. (15) from the induced nucleon decay operator in Eq. (7). The loop integral is dominated by the infrared region, and upon integrating over the loop momentum, the matching relation takes the form,
$\frac{1}{\Lambda_{nn\rightarrow\bar{\nu}\bar{\nu}}^{2}}$ $=$ $\frac{1}{16\pi^{2}}\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}^{2}}\frac{1}{6}\prod_{i=1,2}\Big(\frac{\lambda_{i}v_{S_{i}}m_{\phi_{i}}}{M_{DM_{i}}^{2}}\Big)\sim\frac{1}{16\pi^{2}}\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}^{2}}\frac{1}{6}\Big(\frac{\lambda_{1}v_{S_{1}}\lambda_{2}v_{S_{2}}}{M_{DM}^{2}}\Big)\Big|_{m_{\phi_{i}}\gg\lambda_{i}v_{S_{i}}}$ (16)
$=$ $\frac{1}{16\pi^{2}}\frac{1}{\Lambda_{n\rightarrow\bar{\nu}}^{2}}\frac{1}{6}R_{\bar{\nu}\bar{\nu}}\ ,$
where $M_{DM_{i}}^{2}=\lambda_{i}v_{S_{i}}m_{\phi_{i}}+m_{\phi_{i}}^{2}$ for $i=1,2$, corresponding to $\Phi_{1}$ and $\Phi_{2}$. In the simplifying limit $m_{\phi_{1}}\simeq m_{\phi_{2}}$, we identify $M_{DM_{1}}=M_{DM_{2}}\equiv M_{DM}$. The dimensionless quantity,
$R_{\bar{\nu}\bar{\nu}}=\frac{\lambda_{1}v_{S_{1}}\lambda_{2}v_{S_{2}}}{M_{DM}^{2}}\ll 1\ ,$
encapsulates the suppression associated with the spontaneous breaking of $U(1)_{B}\times U(1)_{L}$ and will be treated as a free parameter in the subsequent analysis.
From the effective Lagrangian, the decay width corresponding to the di-nucleon decay channel $nn\to\bar{\nu}\bar{\nu}$, takes the form,
$\Gamma_{nn\to\bar{\nu}\bar{\nu}}\sim\frac{m_{n}^{5}}{\Lambda_{nn\to\bar{\nu}\bar{\nu}}^{4}}\ ,$ (17)
where $m_{n}$ is the neutron mass. And the present experimental lower limit on the di-nucleon lifetime, quoted in Table (1), translates into the bound,
$\frac{1}{\Lambda_{nn\to\bar{\nu}\bar{\nu}}^{2}}<1.43\times 10^{-31}\ \mathsf{GeV}^{-2}.$ (18)
In addition, by combining this result with the loop-induced matching relation discussed earlier, one can extract an independent constraint on the induced decay operator. For a benchmark choice of dark matter mass $M_{DM}=2\ \mathsf{GeV}$ and loop suppression factor $R_{\bar{\nu}\bar{\nu}}\sim 10^{-7}$, the corresponding lower bound on the effective scale can be estimated as,
$\Lambda_{n\to\bar{\nu}}>4.2\times 10^{7}\ \mathsf{GeV}.$ (19)
This demonstrates that di-nucleon searches, despite being loop-suppressed, can still provide a meaningful and complementary probe of the underlying baryon-number-violating dynamics.
@@
We now consider a Standard Model singlet dark sector scalar field $S$ carrying non-zero baryon number with interactions described by,
$\mathcal{L}_{int}=\Big(\lambda_{1}m_{\phi}S\Phi^{2}+\lambda_{2}S^{\dagger}\Psi^{T}\Psi+c.c\Big)+m_{\phi}^{2}\Phi^{\dagger}\Phi+m_{\psi}\bar{\Psi}\Psi\ .$ (20)
Analogous to the previous case, di-nucleon decay $nn\to\pi^{0}\pi^{0}$ arises at one-loop from the neutron decay operators in Eq. (10), as illustrated in Fig. (4) after the spontaneous breaking of $U(1)_{B}$ symmetry via the vacuum expectation values $v_{S}=\langle S\rangle$. In the unitary gauge, the gauge boson associated with the $U(1)_{B}$ symmetry "eats-up" the Goldstone boson to become heavy.
@@
The hadronic effective Lagrangian governing the di-neutron decay channel $nn\to\pi^{0}\pi^{0}$ can be expressed in terms of the baryon number violating four-fermion operator,